# ASE: Practical Acoustic Speed Estimation Beyond Doppler via Sound Diffusion Field

### v1

@@ 3 Acoustic Diffusion Speed Model
Conventional approaches for speed estimation relying on DFS can only measure the radical speed, making it a location- and direction-dependent solution, as shown in Fig. 1. In this section, we will introduce a novel theoretical model for capturing the complete speed. Our model is primarily inspired by room acoustics [26]. We first briefly review statistical room acoustics and then establish the model that works with channel measurements on commodity audio devices.
3.1 Limitations of DFS
@@
Existing DFS-based acoustic speed estimation implicitly or explicitly relies on a simplistic propagation model. These methods typically involve identifying the primary reflection path between the human body and audio devices and deducing the DFS from the corresponding bins. Consequently, these techniques yield only partial information regarding speed estimation [86, 41, 89, 31], specifically the radial speed $v_{r}$ between the human and acoustic device pair, neglecting the tangent component $v_{t}$, as illustrated in Fig. 1. Essentially, if we only consider one or a few reflection paths, the tangent component will be inevitably lost. Fortunately, we notice that the acoustic signals will be scattered by the numerous reflectors to different directions in the environment, as can be seen from Fig. 3(b) and Fig. 4. If we can efficiently leverage the multi-path reflections, it equivalently creates various estimations of speed along different directions. Conceptually, the reflectors can be regarded as many ”virtual speakers”, offering numerous speed observations of different tangent components. From a statistical perspective, we can imagine these speed components are synthesized into a complete estimation of speed $v$, which contains both racial speed $v_{r}$ and tangent speed $v_{t}$. To this end, we explore the propagation properties of sound and derive the speed information from a statistical approach. We will first introduce acoustic diffusion model in §3.2 and generalize it to CSI in §3.3.
3.2 Sound Diffusion Field
@@
Given a bounded sound propagation scenario (e.g., indoor space), the sound energy is partly reflected/scattered by the obstacles, as shown in Fig. 3. Considering a moving target in space that continuously distorts the sound propagation, a diffuse sound field will be created due to the continuous redistribution of sound energy, especially in an environment with rich diffuseness such as a room. Thus according to the acoustic wave equation [44], we can express the sound pressure as $\mathbf{p}=p(x,t)=Ae^{j(\omega t-\vec{k}\cdot\vec{r})}$ , where $A$ denotes the amplitude, $\omega$ is the frequency of the wave, $\vec{k}$ indicates the propagation vector, and $c$ is the propagation speed of sound. $r$ is the position vector, i.e., $\vec{r}=x\vec{i}+y\vec{j}+z\vec{k}$. Assume the space is excited by a limited-band signal, and consider two adjacent observation points $a$ and $b$ separated by a distance of $x$, as shown in Fig. 3. The sound pressure is expressed as $p_{a}(t)=A\cos(\omega t-\phi)$ and $p_{b}(t)=A\cos\left(\omega t-\phi-kx\cos(\theta)\right)$, respectively, where $\theta$ denotes the direction of the incident sound wave, $A$ and $\phi$ are random amplitudes and angles, and $k$ and $\omega$ represent the wave number and center frequency of the sound signals. The correlation coefficient of the sound energy over space provides insight into the degree of diffuseness. Specifically, the spatial correlation over $p_{a}(t)$ and $p_{b}(t)$ is expressed as
$\psi_{p}=\dfrac{\overline{p_{a}(t)p_{b}(t)}}{\sqrt{\overline{p_{a}(t)^{2}}\cdot\overline{p_{b}(t)^{2}}}},$ (3)
@@
$\psi_{p}\left(x,\theta\right)=\cos\left(kx\cos\theta\right).$ (4)
[…]
Further considering a target moving from point $a$ to $b$ at a speed of $v$, we have $x=v\tau$, where $\tau$ is the traveling time. Thus, the above equation can be written as
$\psi_{p}(v;\tau)=\frac{\sin(kv\tau)}{kv\tau}.$ (7)
It bridges the spatial correlation of the sound pressure $\psi_{p}$ with a target’s moving speed $v$, independent of the moving direction and location. Note that in both scenarios, we average diffuseness across all directions, which differs from DFS that only takes the radial speed into consideration, promising a potential approach for speed estimation beyond the Doppler-based approach. To achieve ASE, however, we need to investigate how to obtain $\psi_{p}$ on commodity audio devices.
3.3 Acoustic Speed Estimation
@@
Figure 5: ACF curves for two different speeds and the theoretical function for $v_{0}=1m/s$.
Channel Power: Indoor space proves to be an environment with rich diffuseness [26], where the above acoustic diffusion model applies. However, it is infeasible to directly measure the sound pressure without professional equipments. It is crucial to efficiently estimate sound pressure on common devices. In ASE, we approximate the sound pressure as the power of the sound diffusion field. Sound pressure can be considered as the equivalent of the power of CSI [27]. CSI is the linear aggregation of multi-path components, which can then be decomposed into two parts, i.e., the static part $\Lambda_{S}$ attributed to static scatterers and the dynamic part $\Lambda_{D}$ contributed by dynamic scatterers, typically from human subjects, as shown in Fig. 4. To sum up, we can model the relationship between channel power and sound pressure as
$G(f,t)$ $=|H(f,t)|^{2}+\epsilon(f,t)\approx|p(f,t)|^{2}+n(f,t)$ (8)
@@
where $H(f,t)$ represents CSI at time $t$ and subcarrier $f$. Practically, the static components are cancelled by removing the time average from $H(f,t)$. $\epsilon(t,f)$ and $n(t,f)$ are noise terms, with a variance of $\Sigma_{N}(f)$. $p(f,t)$ is the sound pressure, which is further decomposed as the dynamic and static components. $p_{i}(f,t)$ is the sound pressure contributed by the $i^{\mathrm{th}}$ scatterer, which can also be deemed as mutually independent of $p_{j}(f,t)$, $\forall i\neq j$, in sound diffusion field. Next, we model the spatial properties of $G(f,t)$ for speed estimation.
@@
$\psi_{G}(f,t,\tau)=\frac{\mathbb{E}\left[G(f,t)G^{H}(f,t+\tau)\right]}{\mathbb{E}\left[G(f,t)G^{H}(f,t)\right]}\triangleq\frac{\mathcal{R}_{1}}{\mathcal{R}_{2}}.$ (9)
@@
$\mathcal{R}_{1}$ $\approx\mathbb{E}\left[\bm{P_{D}}(t)\bm{P_{D}}^{H}(t+\tau)\right]+\mathbb{E}\left[\bm{N}(t)\bm{N}^{H}(t+\tau)\right]$ (10)
@@
$\mathcal{R}_{2}=\bm{\Theta_{D}^{H}}\cdot\mathbf{I_{N_{D}}}+\Sigma_{N}(f).$ (11)
@@
Considering a single moving target, we can assume that all the dynamic scatterers contributed by the target share approximately the same speed $v_{i}$ as the walking speed $v$. The assumption holds in practice as for human targets, the major reflection energy from the torso dominates that from limbs. Notably, prior DFS-based work implicitly adopted similar assumptions, e.g., hand gesture works usually neglect body motions [89, 85, 30]. To this end, we get $\bm{\Psi_{D}}|_{\tau=\tau_{j}}=\psi_{p}(v;\tau_{j})$. With this, we can compute Eq. (9) as
$\psi_{G}(f,\tau)=\bm{\widehat{\Theta_{D}^{H}}}\cdot\left.\bm{\Psi_{D}}\right|_{\tau}=\bm{\widehat{\Theta_{D}^{H}}}\cdot\psi_{p}(v;\tau),$ (13)
where $\bm{\widehat{\Theta_{D}^{H}}}$ is the broadcast aggregation of $\bm{\Theta_{D}^{H}}$ and $\Sigma_{N}(f)$. With the above Eq. (13), we bridge the ACF of the CSI and that of sound pressure as a function of speed, for the first time, offering a completely different approach to acoustic speed estimation by calculating the ACF of the CSI measured on commodity audio devices. Practically, with the CSI time series as input, we can use the sample ACF $\tilde{{\psi}_{G}}(f,\tau)$, where a noise term $\mu(f,\tau)$ will be added to ${\psi}_{G}(f,\tau)$.
@@
$\begin{gathered}\min_{v}\left|x_{0}-\frac{\lambda(f)}{2\pi}v\tau_{s}\right|,\\ \begin{aligned} &\text{s.t.}\quad x_{0}=\min_{x}\left\{x:\frac{d\psi_{p}(x)}{dx}=0\land\frac{d^{2}\psi_{p}(x)}{dx^{2}}<0\right\},\end{aligned}\end{gathered}$ (14)
@@
Figure 6: Transmitter and Receiver Design of ASE. Tx: Band modulation (§4.1) and I/Q modulation (§4.2); Rx: Receiver Demodulation (§4.2) and Channel Estimation (§4.1)
4 ASE Design
@@
4.1 Baseband Sequence Selection
@@
Transmitted Sequence: CSI is by default not available in acoustic sensing, and we send certain signals to probe the channel. Normally, there are three main types of waveform in acoustic sensing [9]: pure-tone, Pseudo-Noise (PN) code, and FMCW. While FMCW signals are widely used in acoustic sensing [12, 29], they are not standard impulse signals and thus introduce CIR estimation errors. As for the sole impulse signal, the energy of the short-time signal would fade out quickly. Therefore, we seek the PN sequence for CIR estimation in ASE. PN sequence consists of equally-spaced Dirac impulses, signs of which alternate with specific rules. It is a noise-like signal with statistical randomness. We choose Kasami sequence [25] as our sensing waveform, mainly due to its superior orthogonality and tolerance to interference. Particularly, the mutual orthogonality of Kasami sequences is critical to our OTDM design, as detailed in §2.
@@
Modulation to Inaudible Band: ASE uses only the inaudible sounds for sensing and modulates the sensing signals on the acoustic band of 17 kHz to 24 kHz, the pseudo-ultrasonic band supported by most commodity devices today. Since the PN code including Kasami has a spreading spectrum over the whole band, we should modulate the Kasami sequence to our desired band. The modulation process is shown in Figure 6. The conversion of the full-band signal into a band signal is achieved by either temporal [85] or spectral [55] interpolation. In ASE, we use the frequency domain interpolation. We perform $N$-point FFT to obtain the frequency domain information, where $N$ is the original length of the sequence. Zero Padding is performed between positive and negative frequency, creating a new sequence of length $N_{s}$. Then we perform IFFT to obtain the temporal signal. The band of the interpolated signal is thus transformed to $\frac{N}{N_{s}}f_{s}$. For a single-sequence signal, we can multiply it with the carrier signal to move the frequency band towards central frequency $f_{c}$. In ASE, we choose $f_{c}=20.25$kHz. We use 63-point Kasami sequences and modulate them to $N_{s}=512$, achieving a bandwidth of 5.9kHz. The transmission signal is coded via PCM and decoded at the receiver’s end.
@@
Figure 7: Illustration of Motion Detection. (a) ACF w/ and w/o movement. ACF without motion has considerably lower ZC counts. (b) Extracted ZCC feature to detect motion.
4.2 Transmission Scheme

### v2

@@ 3. Acoustic Diffusion Speed Model
Conventional approaches for speed estimation relying on DFS can only measure the radial speed, making it a location- and direction-dependent solution, as shown in Fig. 1. In this section, we will introduce a novel theoretical model for capturing the complete speed. Our model is primarily inspired by room acoustics (Kuttruff, 2000). Below, we first briefly review statistical room acoustics and then establish the model that works with channel measurements on commodity audio devices.
3.1. Limitations of DFS
@@
Existing DFS-based acoustic speed estimation implicitly or explicitly relies on a simplistic propagation model. These methods typically involve identifying the primary reflection path between the human body and audio devices and deducing the DFS from the corresponding bins. Consequently, these techniques yield only partial information regarding speed estimation (Zhang et al., 2018; Niu et al., 2022), specifically the radial speed $v_{r}$ between the human and acoustic device pair, neglecting the tangent component $v_{t}$, as illustrated in Fig. 1. Essentially, if we only consider one or a few reflection paths, the tangent component will inevitably be lost. Fortunately, we notice that the acoustic signals will be scattered by the numerous reflectors to different directions in the environment, as can be seen from Fig. 3(b) and Fig. 4. If we can efficiently leverage the multi-path reflections, it equivalently creates various estimations of speed along different directions. Conceptually, the reflectors can be regarded as many ”virtual speakers”, offering numerous speed observations of different tangent components. From a statistical perspective, we can imagine these speed components are synthesized into a complete estimation of speed $v$, which contains both racial speed $v_{r}$ and tangent speed $v_{t}$. To this end, we explore the propagation properties of sound and derive the speed information from a statistical approach. We will first introduce the acoustic diffusion model in §3.2 and generalize it to CSI in §3.3.
3.2. Sound Diffusion Field
Figure 4. Diffusive Scattering in the room. (a) Illustration of diffusion scattering and specular reflections. The acoustic waves not only experience specular reflections, but also diffusive scatterings on any non-flat planes. (b) A simulation result of how an acoustic wave is reflected and scattered in a common room. The black point emits the acoustic rays with a frequency range of 17 kHz- 20 kHz. By employing the rich multi-path semantics, these scatterers are creating ”virtual speakers”, thus creating different views of speed estimation.
@@ 3.2. Sound Diffusion Field
Given a bounded sound propagation scenario (e.g., indoor space), the sound energy is partly reflected/scattered by the obstacles, as shown in Fig. 3. Considering a moving target in space that continuously distorts the sound propagation, a diffuse sound field will be created due to the continuous redistribution of sound energy, especially in an environment with rich diffuseness such as a room. Thus, according to the acoustic wave equation (Pierce, 2019), we can express the sound pressure as $\mathbf{p}=p(x,t)=Ae^{j(\omega t-\vec{k}\cdot\vec{r})}$ , where $A$ denotes the amplitude, $\omega$ is the frequency of the wave, $\vec{k}$ indicates the propagation vector, and $c$ is the propagation speed of sound. $r$ is the position vector, i.e., $\vec{r}=x\vec{i}+y\vec{j}+z\vec{k}$. Assume the space is excited by a limited-band signal, and consider two adjacent observation points $a$ and $b$ separated by a distance of $x$, as shown in Fig. 3. The sound pressure is expressed as $p_{a}(t)=A\cos(\omega t-\phi)$ and $p_{b}(t)=A\cos\left(\omega t-\phi-kx\cos(\theta)\right)$, respectively, where $\theta$ denotes the direction of the incident sound wave, $A$ and $\phi$ are random amplitudes and angles, and $k$ and $\omega$ represent the wave number and center frequency of the sound signals. The correlation coefficient of the sound energy over space provides insight into the degree of diffuseness. Specifically, the spatial correlation over $p_{a}(t)$ and $p_{b}(t)$ is expressed as
(3) $\psi_{p}=\dfrac{\overline{p_{a}(t)p_{b}(t)}}{\sqrt{\overline{p_{a}(t)^{2}}\cdot\overline{p_{b}(t)^{2}}}},$
where $\psi_{p}$ represents the correlation coefficient, and $\overline{\left.\cdot\right.}$ denotes operation of time average. Since sound is a plane wave, we can derive the time average of $p_{a}(t)^{2}$ and $p_{b}(t)^{2}$ both as $\frac{A^{2}}{2}$, and meanwhile, the average of the product of sound pressure is computed as $\overline{p_{a}(t)p_{b}(t)}=\frac{1}{\Delta t}\int_{t}^{t+\Delta t}p_{a}(t)p_{b}(t)dt=\frac{A^{2}}{2}\cos(kx\cos\theta)$. Therefore, by substituting them into Eq. (3), we derive the correlation for the direction of incidence $\theta$ as
(4) $\psi_{p}\left(x,\theta\right)=\cos\left(kx\cos\theta\right).$
(5) $\psi_{p}\left(x\right)=\frac{1}{2\pi}\int_{0}^{2\pi}\psi\left(x,\theta\right)d\theta=J_{0}(kx),$
where $J_{0}(x)=\frac{1}{2\pi}\int_{0}^{2\pi}\cos(x\cos\theta)d\theta$ is the $0^{\mathrm{th}}$-order Bessel function of its first kind. If we consider 3D scattering model, i.e., the sound waves are distributed in a sphere, we can acquire the correlation coefficient as,
(6) $\psi_{p}\left(x\right)=\frac{1}{4\pi}\int_{0}^{\pi}\int_{0}^{2\pi}\psi\left(x,\theta\right)d\phi d\theta=\frac{\sin(kx)}{kx},$
@@
(7) $\psi_{p}(v;\tau)=\frac{\sin(kv\tau)}{kv\tau}.$
It bridges the spatial correlation of the sound pressure $\psi_{p}$ with a target’s moving speed $v$, independent of the moving direction and location. Note that in both scenarios, we average diffuseness across all directions, in contrast to DFS, that only takes the radial speed into consideration, promising a potential approach for speed estimation beyond the Doppler-based approach.
Table 1. Examples of different scattering coefficients indoor (Faiz et al., 2012; Training, 2015)
Room Components Scattering Coefficient
Cabinet 010.2
4 Chairs 010.45
1 table + 3 chairs + PC 010.45
Audience 010.7
Carpets 010.3
Irregular Books 010.5
Understanding Sound Diffuseness in Room Environment: Indoor environments inherently exhibit scattering effects. As established by modern room acoustic studies (Faiz et al., 2012; Training, 2015; Kuttruff, 2000), the level of scattering increases with increasing sound frequency and the roughness of surfaces. In common room settings, these scattering effects are ubiquitous and pervasive. Curved walls, textured surfaces, or objects with uneven shapes do not reflect sound in a single, coherent direction. Instead, they disperse sound energy across a wide range of angles, much like how a crumpled piece of paper scatters light compared to a smooth mirror. This significantly increases the scattering coefficient, meaning more sound energy is diffused rather than specularly reflected. As summarized in Tab. 1, typical scattering coefficients for indoor materials range from 0 (no scattering) to 1 (full scattering). Ordinary objects such as chairs and tables act as scatterers, diffusing incident sound energy throughout the space (see Fig. 3(b)). The scattering becomes more pervasive in complex room settings and with irregular reflectors. This implies that modeling sound diffusion in a room offers a new perspective on its acoustic behavior. By adopting a statistical perspective on the sound field, we can analyze the acoustic environment without considering the specific material types of objects, whether they are on Line of Sight (LoS) or Non-Line of Sight (NLoS) paths. In other words, this statistical approach enables us to estimate overall acoustic properties in a more comprehensive way than the DFS method and simplifies the processing by eliminating the need to separate specific paths from the received signal.
To achieve ASE, however, we need to investigate how to obtain $\psi_{p}$ on commodity audio devices.
3.3. Acoustic Speed Estimation
@@
Figure 5. ACF curves for two different speeds and the theoretical function for $v_{0}=1m/s$.
Channel Power: Indoor space proves to be an environment with rich diffuseness (Kuttruff, 2000), where the above acoustic diffusion model applies. However, it is infeasible to directly measure the sound pressure without professional equipments. It is crucial to efficiently estimate sound pressure on common devices. In ASE, we approximate the sound pressure as the power of the sound diffusion field. Sound pressure can be considered as the equivalent of the power of CSI (Leighton and Sands, 1965). CSI is the linear aggregation of multi-path components, which can then be decomposed into two parts, i.e., the static part $\Lambda_{S}$ attributed to static scatterers and the dynamic part $\Lambda_{D}$ contributed by dynamic scatterers, typically from human subjects, as shown in Fig. 4. To sum up, we can model the relationship between channel power and sound pressure as
(8) $G(f,t)$ $=|H(f,t)|^{2}+\epsilon(f,t)\approx|p(f,t)|^{2}+n(f,t)$
@@
where $H(f,t)$ represents CSI at time $t$ and subcarrier $f$. Practically, the static components, including the direct path between the speaker and microphone, are cancelled by removing the time average from $H(f,t)$. $\epsilon(t,f)$ and $n(t,f)$ are noise terms, with a variance of $\Sigma_{N}(f)$. $p(f,t)$ is the sound pressure, which is further decomposed as the dynamic and static components. $p_{i}(f,t)$ is the sound pressure contributed by the $i^{\mathrm{th}}$ scatterer, which can also be deemed as mutually independent of $p_{j}(f,t)$, $\forall i\neq j$, in sound diffusion field. Next, we model the spatial properties of $G(f,t)$ for speed estimation.
@@
(9) $\psi_{G}(f,t,\tau)=\frac{\mathbb{E}\left[G(f,t)G^{H}(f,t+\tau)\right]}{\mathbb{E}\left[G(f,t)G^{H}(f,t)\right]}\triangleq\frac{\mathcal{R}_{1}}{\mathcal{R}_{2}}.$
@@
(10) $\mathcal{R}_{1}$ $\approx\mathbb{E}\left[\bm{P_{D}}(t)\bm{P_{D}}^{H}(t+\tau)\right]+\mathbb{E}\left[\bm{N}(t)\bm{N}^{H}(t+\tau)\right]$
@@
(11) $\mathcal{R}_{2}=\bm{\Theta_{D}^{H}}\cdot\mathbf{I_{N_{D}}}+\Sigma_{N}(f).$
@@
Considering a single moving target, we can assume that all the dynamic scatterers contributed by the target share approximately the same speed $v_{i}$ as the walking speed $v$. The assumption holds in practice as for human targets, the major reflection energy from the torso dominates that from limbs. Notably, prior DFS-based work implicitly adopted similar assumptions, e.g., hand gesture works usually neglect body motions (Zhang et al., 2023b; Yun et al., 2017; Li et al., 2022c). To this end, we get $\bm{\Psi_{D}}|_{\tau=\tau_{j}}=\psi_{p}(v;\tau_{j})$. With this, we can compute Eq. (9) as
(13) $\psi_{G}(f,\tau)=\bm{\widehat{\Theta_{D}^{H}}}\cdot\left.\bm{\Psi_{D}}\right|_{\tau}=\bm{\widehat{\Theta_{D}^{H}}}\cdot\psi_{p}(v;\tau),$
where $\bm{\widehat{\Theta_{D}^{H}}}$ is the broadcast aggregation of $\bm{\Theta_{D}^{H}}$ and $\Sigma_{N}(f)$.
With the above Eq. (13), we bridge the ACF of the CSI and that of sound pressure as a function of speed, for the first time, offering a completely different approach to acoustic speed estimation by calculating the ACF of the CSI measured on commodity audio devices. Practically, with the CSI time series as input, we can use the sample ACF $\tilde{{\psi}_{G}}(f,\tau)$, where a noise term $\mu(f,\tau)$ will be added to ${\psi}_{G}(f,\tau)$.
@@
[…]
@@
Figure 6. Transmitter and Receiver Design of ASE. Tx: Band modulation (§4.1) and I/Q modulation (§4.2); Rx: Receiver Demodulation (§4.2) and Channel Estimation (§4.1).
Multipath Effect: The previous practice separates different paths from the CSI for modeling the multipath propagation. The multipath effect can be harmful for those approaches, as the multipath components will be hard to extract in complicated scenarios. However, our sound diffusion model does not involve decomposing each multipath, instead aggregates the speed information from all the multipaths. Intuitively, we analyze the acoustic channel from a statistical perspective, where the multipath information is aggregated together for speed estimation. In other words, we leverage the diffusiveness of the room to infer the speed. To this end, it is practical for realistic room environments and robust to multipath effects. At the same time, we do not need to focus on the specific material properties within the room, as our method integrates information holistically, ensuring that no single material acts as a bottleneck.
@@
4. ASE Design
@@
4.1. Baseband Sequence Selection
@@
Transmitted Sequence: CSI is by default not available in acoustic sensing, and we send certain signals to probe the channel. Normally, there are three main types of waveform in acoustic sensing (Cai et al., 2022): pure-tone, Pseudo-Noise (PN) code, and FMCW. While FMCW signals are widely used in acoustic sensing (Cheng and Lou, 2021; Li et al., 2022b), they are not standard impulse signals and thus introduce CIR estimation errors (Mao et al., 2020; Viikari et al., 2008). As for the sole impulse signal, the energy of the short-time signal would fade out quickly. Therefore, we seek the PN sequence for CIR estimation in ASE. PN sequence consists of equally spaced Dirac impulses, the signs of which alternate with specific rules. It is a noise-like signal with statistical randomness. We choose the Kasami sequence (Kasami, 1966) as our sensing waveform, mainly due to its superior orthogonality and tolerance to interference. Particularly, the mutual orthogonality of Kasami sequences is critical to our OTDM design, as detailed in §2.
@@
Modulation to Inaudible Band: ASE uses only the inaudible sounds for sensing and modulates the sensing signals on the acoustic band of 17 kHz to 24 kHz, the pseudo-ultrasonic band supported by most commodity devices today. Since the PN code, including Kasami, has a spreading spectrum over the whole band, we should modulate the Kasami sequence to our desired band. The modulation process is shown in Figure 6. The conversion of the full-band signal into a band signal is achieved by either temporal (Yun et al., 2017) or spectral (Sun et al., 2018) interpolation. In ASE, we use the frequency domain interpolation. We perform $N$-point FFT to obtain the frequency domain information, where $N$ is the original length of the sequence. Zero Padding is performed between positive and negative frequencies, creating a new sequence of length $N_{s}$. Then we perform IFFT to obtain the temporal signal. The band of the interpolated signal is thus transformed to $\frac{N}{N_{s}}f_{s}$. For a single-sequence signal, we can multiply it by the carrier signal to move the frequency band towards the central frequency $f_{c}$. In ASE, we choose $f_{c}=20.25$kHz. We use 63-point Kasami sequences and modulate them to $N_{s}=512$, achieving a bandwidth of 5.9kHz. The transmission signal is coded via PCM and decoded at the receiver’s end.
@@
Figure 7. Illustration of Motion Detection. (a) ACF w/ and w/o movement. ACF without motion has considerably lower ZC counts. (b) Extracted the ZCC feature to detect motion.
4.2. Transmission Scheme